OpenStudy (anonymous):

Write an equation in slope intercept form of the line through point P(8,2) with slope 4.....A. y-2=4(x-8). B. y-8=4(x-2). C. Y=4x+2. D. Y=4x-30

OpenStudy (anonymous):

slope intercept form : y - y1 = m(x - x1) slope(m) = 4 (8,2)...x1 = 8 and y1 = 2 now we sub y - 2 = 4(x - 8)....thats it....it's that simple